OpenStudy (anonymous):

if cos75=x, what is the value in terms of x: (sin165-cos75)/1+(cos15)(cos165)

OpenStudy (jdoe0001):

\(\bf \cfrac{sin(165^o)-cos(75^o)}{1-cos(15^o)cos(165^o)}\quad ?\)

OpenStudy (anonymous):

for the denominator, it is 1+, not 1- :) so this means you can help me? :D

OpenStudy (jdoe0001):

\(\bf \cfrac{sin(165^o)-cos(75^o)}{1+cos(15^o)cos(165^o)}\) right?

OpenStudy (anonymous):

yeap

OpenStudy (jdoe0001):

\( \bf \cfrac{sin(165^o)-cos(75^o)}{1+cos(15^o)cos(165^o)}\implies \cfrac{[sin(90^o+75)]-cos(75^o)}{1+cos(90^o+75^o)cos(90^o+75^o)}\\ \quad \\ \cfrac{[sin(90^o)cos(75^o)+cos(90^o)sin(75^o)]-cos(75^o)} { \begin{array}{llll} 1+[cos(90^o)cos(75^o)+sin(90^o)sin(75^o)]\\\ \bf [cos(90^o)cos(75^o)-sin(90^o)sin(75^o)] \end{array}}\)

OpenStudy (jdoe0001):

anyhow.... I need to dash keep in mind that \(\bf sin(90^o)=1\qquad cos(90^o)=0\qquad cos(75^o)=x\) replace accordingly :)

OpenStudy (anonymous):

you are blessing on this earth thank you so much!!!!